Instructions: Solve each system by the substitution method.
What is the solution for 3x+y=-7
x-y=-5

Answer by jai_kos(139)

Solution:
Given
3x + y = -7 (1)

x - y = -5 (2)
Consider equation (2), we can rewrite the equation (2) as
x + 5 = y (3)
Substitute the equation (3) in equation (1), we get
3x + (x + 5) = - 7
3x + x + 5 = -7
4x + 5 = -7
4x = - 7 – 5
4x = -12
x = -12 / 4
x = -3 (4)

Put x = -3 in equation (1), we get
3(-3) + y = -7
-9 + y = - 7
y = - 7 + 9
y = 2
Therefore the value of x = -3 and y = 2